Set up the differential equation for transverse waves in a string.
1
Expert's answer
2019-06-11T10:01:41-0400

The equation for transverse wave in a string is:


"\frac{\partial^2 \xi}{\partial t^2}=\frac{T}{\mu}\frac{\partial^2 \xi}{\partial x^2}"

where "\xi" - displacement from equilibrium, "T" - tension in the string, "\mu" - mass of the string per unit length